OpenStudy (anonymous):

Im a freshman taking algebra 2 next year with juniors, and on my summer h.w. i need to do algorithms. i have no clue... HELP!

OpenStudy (anonymous):



OpenStudy (anonymous):

and above it says "suppose that (x in a triangle) means x+5, (x in a square) means x^2, and (x in a circle) means 2x-3."

OpenStudy (anonymous):

is the 4 an x?

OpenStudy (anonymous):

in the triangle it is a four, thats what the diagram says

OpenStudy (anonymous):

x+5 +x^2 +2x-3 =x^2+3x+2=(x+2)(x+1) now put in 4 for x

OpenStudy (anonymous):

so basically your just adding them altogether in an equation, and then substituting for x?

OpenStudy (phi):

x in triangle, with x=4 gives 9 in circle. 9 in circle is 2*9-3 = 15 15 in square is 15*15 = 225

OpenStudy (anonymous):

thank you